OpenStudy (anonymous):

3x+4=5x-6 what is x?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

3x+4=5x-6 5x-3x=4+6 2x=10 x=5

OpenStudy (anonymous):

well maybe you can help me with my next problem?

OpenStudy (aroub):

what is it?

OpenStudy (anonymous):

5x+11+109+180 what is x?

OpenStudy (anonymous):

if tyhe equation is equal to 0 then 5x+300=0 5x=-300 x=-60

OpenStudy (aroub):

do u know how to do it? nikole

OpenStudy (anonymous):

yes kinda

OpenStudy (aroub):

can u ask another one and u do it please

OpenStudy (anonymous):

i got 12.4 for the qusetion i just asked?

OpenStudy (aroub):

hahaha how?

OpenStudy (anonymous):

5x+11+109=180 5x+118=180 5x=62 x=12.4

OpenStudy (anonymous):

there is no -180

OpenStudy (aroub):

it has to be =-180

OpenStudy (aroub):

5x+11+109=-180

OpenStudy (anonymous):

it was 12 i was correct

OpenStudy (aroub):

but u made a very long way just add the numbers and the variables together and u will find it

OpenStudy (aroub):

wait i want to solve it

OpenStudy (aroub):

300+5x=0 5x=-300 x=-60

OpenStudy (anonymous):

where are you getting 300

OpenStudy (aroub):

add 109+180+11=300

OpenStudy (aroub):

try it

OpenStudy (anonymous):

tht wasnt the problem i gave you

OpenStudy (aroub):

yes it is the second one